OpenStudy (anonymous):

1/(1 - sinx) - 1/(1 + sinx) I believe the answer is 1/cos^2x, but I don't know why. Can someone help walk me through the steps?

zepdrix (zepdrix):

\[\large\rm \frac{1}{1-\sin x}-\frac{1}{1+\sin x}\] Do you remember your Pythagorean Identity involving sine and cosine? :) Here is one way we can write that relationship,\[\large\rm 1-\sin^2x=\cos^2x\]

zepdrix (zepdrix):

Notice that the left side of this identity is really the `difference of two squared numbers`\[\large\rm 1^2-\sin^2x\]This will factor into conjugates,\[\large\rm =(1-\sin x)(1+\sin x)\]

zepdrix (zepdrix):

That should give you a sense of what direction we'll be heading in our steps :)

zepdrix (zepdrix):

If we're able to get some (1-sin x)(1+sin x) in our denominators, we'll be able to apply our Pythagorean Identity.

zepdrix (zepdrix):

So if that's the case, I guess we would like to have a (1+sin x) in the bottom and top of our first fraction,\[\large\rm \color{royalblue}{\frac{1+\sin x}{1+\sin x}}\cdot\frac{1}{1-\sin x}-\frac{1}{1+\sin x}\]

zepdrix (zepdrix):

Which becomes,\[\large\rm \frac{1+\sin x}{(1+\sin x)(1-\sin x)}-\frac{1}{1+\sin x}\]

zepdrix (zepdrix):

What do you think gates? Too confusing?

OpenStudy (anonymous):

It's fairly confusing. That's probably why I didn't quite absorb the information in class. This whole chapter is crazy.

OpenStudy (anonymous):

\[\frac{1}{1-b}-\frac{1}{1+b}\] \[=\frac{1+b-(1-b)}{1-b^2}=\frac{2}{1-b^2}\] it is 90% algebra

zepdrix (zepdrix):

Well you're familiar with the idea of a `common denominator`, yes?\[\large\rm \frac{1}{a}-\frac{1}{b}\quad=\quad \frac{b}{ab}-\frac{a}{ab}\quad=\frac{b-a}{ab}\]That's a good starting point for this problem. We would like a new denominator which has both denominators as factors. So our new denominator should consist of both (1-sin x) and (1+sin x). So we go from here,\[\large\rm \frac{1}{(1-\sin x)}-\frac{1}{(1+\sin x)}\]to here,\[\large\rm \frac{\color{royalblue}{(1+\sin x)}}{\color{royalblue}{(1+\sin x)}(1-\sin x)}-\frac{\color{orangered}{(1-\sin x)}}{(1+\sin x)\color{orangered}{(1-\sin x)}}\]

OpenStudy (anonymous):

Yeah so this all leads back down to 1/1-sinx - 1/1+sin, but where do I go from there?

zepdrix (zepdrix):

Well we don't want to go backwards. We're turning it into this big mess for a specific reason. Now we're able to combine it into a single fraction,\[\large\rm \frac{(1+\sin x)-(1-\sin x)}{(1+\sin x)(1-\sin x)}\]

zepdrix (zepdrix):

If we drop the brackets in the numerator, distributing the negative,\[\large\rm \frac{1+\sin x-1+\sin x}{(1+\sin x)(1-\sin x)}\]We can combine like-terms. Looks like the 1's will cancel out, ya?

OpenStudy (anonymous):

Yup

zepdrix (zepdrix):

\[\large\rm \frac{2\sin x}{(1+\sin x)(1-\sin x)}\]So that simplifies down the top as far as we're able to go.

zepdrix (zepdrix):

And again, bottom is an identity. If you don't quite understand it, maybe just "believe it" for right now.\[\large\rm (1+\sin x)(1-\sin x)=\cos^2x\]

zepdrix (zepdrix):

\[\large\rm \frac{2\sin x}{\cos^2x}\]

OpenStudy (anonymous):

Oh okay, thanks so much!

zepdrix (zepdrix):

np